# Non-linear Welfare-Aware Strategic Learning

### v1

@@ 3.2 Welfare of different parties
When $u(D(x),y)=\textbf{1}(D(x)=y)$, decision-maker welfare defined above reduces to the prediction accuracy $P(D(X)=Y)$. In Sec. 5 and 6, we focus on the special case with $\text{DW}(f)=\operatorname{\mathbb{E}}_{x,y\sim P_{XY}}[\textbf{1}(D(x)=y)]$, but the results and analysis can be extended to other utility functions.
@@
It measures the increase in the overall population qualification caused by the agent best response. This notion has been theoretically and empirically studied under linear settings, e.g., (Kleinberg and Raghavan, 2020; Bechavod et al., 2022).
@@
It illustrates whether there are agents suffering from deteriorating qualifications after they best respond. This notion has been studied in (Bechavod et al., 2022) and was interpreted as "do no harm" constraint, i.e., the deployed decisions should avoid adversely impacting agents. Rosenfeld et al. (2020) has shown that agent safety can not be automatically guaranteed even under linear settings, and they proposed an algorithm to ensure safety in general settings. However, they did not consider agent improvement which we believe is an equally important measure of social welfare.
@@
Previous literature rarely discussed the welfare of agents. The closest notion was social burden (Milli et al., 2019) defined as the minimum cost a positively-labeled user must pay to obtain a positive decision, where the authors assume agents can only “game" the decision-maker without improving their true qualifications. Since the agents can always strive to improve in our setting, their efforts are not necessarily a “burden". Thus, we propose a definition of agent welfare which measures the extent to which the agent qualification is under-estimated by the decision-maker, i.e.,
@@ 4 Welfare Analysis
This section explores the relationships between the decision-maker, social, and agent welfare introduced in Sec. 3. Among all prior works, Bechavod et al. (2022) provides a theoretical analysis of social and decision-maker welfare under linear settings. In this section, we first revisit the results in (Bechavod et al., 2022) and show that they may not be applicable to non-linear settings. Then, we examine the relationships between three types of welfare under general settings.
@@ 4.1 Beyond the linear setting
Consider a hiring scenario where a company needs to assess each applicant’s fitness for a position. Each applicant has a feature $X\in[0,1]$ demonstrating his/her ability (e.g., coding ability for a junior software engineer position). Suppose the labeling function $h(x)=-4x\cdot(x-1)$ is quadratic with maximum taken at $x=0.5$, which implies that an applicant’s fitness will decrease if the ability is too low ("underqualified") or too high ("overqualified")111”Overqualified” means an applicant has skills significantly exceeding the requirement. In this case, he/she may fit more senior positions or desire a much higher salary than this position can offer.. Under the quadratic labeling function, if the decision-maker can only select a decision policy $f$ from a linear hypothesis class $\mathcal{F}$ (i.e., violating the requirement that $h\in\mathcal{F}$ in Thm. 4.1), then the optimal policies maximizing decision-maker welfare and agent improvement are completely different, i.e., $\text{DW}(f)$ and $\text{IMP}(f)$ cannot be maximized simultaneously. Moreover, agent safety $\text{SF}(f)$ is also not guaranteed. To illustrate this, consider the example in Fig. 1 with five agents: four green agents with features below $0.5$ and a black one above $0.5$, then the policy that maximizes $\text{IMP}(f)$ can cause the qualification of black agent to deteriorate and lead to negative agent safety $\text{SF}(f)<0$.
@@
Under the same setting of Example 4.2, instead of restricting $f$ to be linear, let the decision-maker deploy $f(x)=h(x)=-4x\cdot(x-1)$. Obviously, $f$ maximizes both the decision-maker welfare and agent welfare. However, it can hurt social welfare. Consider an example shown in Fig. 2 where all agents have the same feature $x=0.4$ and information level $K=1$ (violating the information level requirement of Thm. 4.1). Their cost function is $c(x,x^{\prime})=4\cdot(x-x^{\prime})^{2}$. Since $\nabla f(x)=-8x+4=3.68$, agents derive $Q^{f}_{x}(x^{\prime})=f(x)+3.68\cdot(x^{\prime}-x)=3.68\cdot x^{\prime}-0.512$. Then agent post-response feature will be $x^{*}=x+\frac{3.68}{8}=0.86$. As a result, the agent qualification becomes $h(x^{*})=h(0.86)=0.4816<h(0.4)=0.96$, which is worse than the initial qualification, i.e., agent safety is not optimal by deploying $f=h$.
Figure 2: Violation of agent safety $\text{SF}(f)$ when $f=h$ is quadratic (the blue curve): All agents have the same feature value concentrating on $X=0.4$ (the solid black circle). If applying the ground truth as the decision policy, all agents will respond by moving to $x^{*}=0.86$ (the hollow black circle). Although the policy maximizes $\text{AW}(f)$ and $\text{DW}(f)$, all agents become less qualified after they respond, harming the social welfare. Both $\text{IMP}(f)$ and $\text{SF}(f)$ are smaller than 0.
@@ 4.2 Compatibility analysis of each welfare pair
4.2.1 Decision-maker welfare & agent welfare
@@
4.2.2 Decision-maker welfare & social welfare
Unlike compatibility results in Sec 4.2.1, the decision-maker welfare and either element of social welfare (i.e., agent improvement $\text{IMP}(f)$ or agent safety $\text{SF}(f)$) in general are not compatible, and their compatibility needs all conditions in Thm. 4.1 to be satisfied.
@@ 4.2.2 Decision-maker welfare & social welfare
4.2.3 Agent improvement & agent safety
@@
Suppose the agents have information level $K$ and respond to $f$ based on Def. 3.2. Let $^{*}\in\arg\max_{f\in\mathcal{F}}\text{IMP(f)}$(f) be any policy that maximizes the agent improvement, and let $Q^{f^{*}}_{x}$ be the estimation of $f^{*}$ for agents with feature vector $x$. Then any $f^{*}$ is guaranteed to be safe (i.e., $\text{SF}(f^{*})=0$) regardless of the feature distribution $P_{X}$ and the cost function $c(x,x^{\prime})$ if and only if either of the following scenarios holds: (i) Thm. 4.1 holds; (ii) For each agent with feature vector $x$, we have: $\frac{\partial h}{\partial\widetilde{x}_{i}}\cdot\frac{\partial Q^{f^{*}}_{x}}{\partial\widetilde{x}_{i}}\geq 0$ holds almost everywhere in $\widetilde{x}=[\widetilde{x}_{1},...,\widetilde{x}_{d}]\in\text{dom}(X)$.
@@
4.2.4 Social welfare & agent welfare
@@
Suppose the agents have information level $K$ and respond to $f$ based on Def. 3.2. Denote $\mathcal{F}^{\prime}$ as $\{f:f\in\mathcal{F}~{}and~{}f(x)\geq h(x),\forall x\in dom(x)\}$ be the set of policies maximizing the agent welfare. Then for any $f^{*}\in\arg\max_{f\in\mathcal{F}}\text{IMP}(f)$, $f^{*}$ is guaranteed to also maximize $\text{AW}(f)$ regardless of the feature distribution $P_{X}$ and the cost function $c(x,x^{\prime})$ if and only if either of the following scenarios holds: (i) Thm. 4.1 holds; (ii) $\exists f^{\prime}\in\mathcal{F}^{\prime}$ and constant $C\in\operatorname{\mathbb{R}}^{+}$ to let $Q^{f^{\prime}}_{x}=Q^{f^{*}}_{x}+C$ hold almost everywhere in $dom(x)$.
Again, scenario (i) in Thm. 4.7 is a sufficient condition, while scenario (ii) means that the decision-maker can find a classifier that induces perfect agent welfare while simultaneously giving the agents the same information as $f^{*}$. Note that $Q^{f^{\prime}}_{x}=Q^{f^{*}}_{x}+C$ does not necessarily imply $f^{\prime}=f^{*}+C$ when agents have a limited information level $K$. For example, when agents have $K=2$, we can have $f^{*}=\text{e}^{x}-x-1$ and $f^{\prime}=\frac{x^{2}}{2}+0.5$, where $f^{*}$ belongs to the exponential function family but $f^{\prime}$ belongs to the quadratic function family.
To conclude, only decision-maker welfare and agent welfare can be maximized at the same time relatively easily, other pairs of welfare need strong conditions to align under the general settings. Even the two notions in social welfare (i.e., agent improvement and safety) can easily contradict each other, and the decision-maker welfare and social welfare need the most restrictive conditions (which is exactly Thm. 4.1) to be aligned. These results highlight the difficulties of balancing the welfare of different parties in general non-linear settings, which further motivates the “irreducible" optimization framework we will introduce in Sec. 5.
@@ 5 Welfare-Aware Optimization
Output: Deployed classifier $f\in\mathcal{F}$
@@
10:  Return $\theta$
Algorithm 1 Strategic Welfare-Aware Optimization (STWF)
@@
The optimization (2) can be solved using gradient-based algorithms, as long as the gradient of losses with respect to the model parameter (denoted as $\theta$) exists. Since $f,h$ are continuous, gradients already exist for $\mathbf{\ell_{DW}}$ and $\mathbf{\ell_{AW}}$. For $\mathbf{\ell_{SWF}}$, because it is a function of post-response feature $X^{*}$, we need gradients to exist for $X^{*}$ as well. If we already know how agents respond to the model (i.e., $X^{*}$ as a function of $f$ illustrated in Def. 3.2), then we only need gradients to exist for $\nabla f(x)$ with respect to $\theta$ (i.e., Hessian of $f$). Thus, the objective (2) can be easily optimized if $\nabla f(x)$ is continuous in $\theta$, which is a relatively mild requirement and has also been used in previous literature (Rosenfeld et al., 2020). The complete algorithm for StraTegic WelFare-Aware optimization (STWF) is shown in Alg. 1.
@@
6:  Train $\Delta_{\phi}$ with $\mathcal{S}$
7:  Return $\phi$
Algorithm 2 AgentResponse
Alg. 2 requires a population for controlled experiments. As pointed out by Miller et al. (2020), we can never understand the agent response without the controlled trials. Though the availability of controlled experiments has been assumed in most literature on performative prediction Perdomo et al. (2020); Izzo et al. (2021), it may be costly to conduct the experiments in practice.
@@ 6 Experiments
(a) Welfare while increasing $\lambda_{1}$ and keeping $\lambda_{2}=0$.
(b) Welfare while increasing $\lambda_{2}$ and keeping $\lambda_{1}=0$.
Figure 3: Effects of adjusting $\lambda_{1},\lambda_{2}$ on the welfare of each party: synthetic data (left plot), German Credit data (middle plot), ACSIncome-CA data (right plot).
We conduct comprehensive experiments on one synthetic and two real-world datasets to validate STWF. We first illustrate the trade-offs for each pair of welfare and examine how they can be balanced by adjusting $\lambda_{1},\lambda_{2}$ in Alg. 1. Then, we compare STWF with previous algorithms that only consider the welfare of a subset of parties. Since many of the existing algorithms also consider fairness across agents from different groups, we also evaluate the unfairness induced by our algorithm in addition to decision-maker, agent, and social welfare. We use a three-layer ReLU neural network to learn labeling function $h$ and logistic regression to train the policy $f$, and $D(x)=\mathbf{1}(x\geq 0.5)$. For each parameter setting, we perform 20 independent runs of the experiments with different random seeds and report the average.
6.1 Datasets
  • ACSIncome-CA data (Ding et al., 2021). It contains income-related data for California residents. Let label $Y$ be whether an agent’s annual income is higher than $50$K USD and let the sensitive attribute $Z$ be sex. We randomly select $20000$ samples from the data of year $2018$. We assume all agents have the same effort budget $c(x,x^{\prime})=\|x^{\prime}-x\|^{2}$.
  • German Credit data (Dua and Graff, 2017). It has $1000$ samples and the label $Y$ is whether the agent’s credit risk is high or low. We adopt the data preprocessed by Jiang and Nachum (2020) where age serves as the sensitive attribute to split agents into two groups. We assume the agents are able to improve the following features: existing account status, credit history, credit amount, saving account, present employment, installment rate, guarantors, residence. We assume all agents have the same effort budget $c(x,x^{\prime})=\|x^{\prime}-x\|^{2}$.
6.2 Balance the Welfare with Different $\lambda_{1},\lambda_{2}$
We first validate STWF and show that adjusting the hyper-parameters $\lambda_{1},\lambda_{2}$ effectively balances the welfare of different parties. The results in Fig. 3 are as expected, where social welfare $\text{SWF}(f)$ increases in $\lambda_{1}$ and agent welfare $\text{AW}(f)$ increases in $\lambda_{2}$. This shows the effectiveness of using our algorithm to adjust the allocation of welfare under different situations. As we analyzed in Sec. 4, under the general non-linear setting, each pair of welfare possibly conflicts with each other. Fig. 3 validates this and it demonstrates the trade-offs between different welfare pairs on two datasets. Specifically, for the synthetic data, increasing social welfare slightly conflicts with decision-maker welfare, while increasing agent welfare drastically conflicts with the decision-maker welfare. But the trade-off between social welfare and agent welfare is not obvious. For the German Credit dataset, increasing social welfare seems to slightly conflict with both agent welfare and decision-maker welfare, while increasing agent welfare has an obvious trade-off with decision-maker welfare. Conversely, for the ACSIncome-CA dataset, increasing social welfare only sacrifices agent welfare, but increasing agent welfare incurs conflicts with both other welfare.
6.3 Comparison with Previous Algorithms
Specifically, we compare STWF with four algorithms:
  • Safety (SAFE) (Rosenfeld et al., 2020): It considers both decision-maker welfare and agent safety. The goal is to train an accurate model without hurting agent qualification.
  • Equal Improvement (EI) (Guldogan et al., 2022): It considers agent improvement by equalizing the probability of unqualified agents becoming qualified from different groups, as opposed to maximizing agent improvement considered in our work.
  • Bounded Effort (BE) (Heidari et al., 2019): It considers agent improvement by equalizing the proportion of new agents becoming qualified after the best response. The difference between BE and EI is a bit subtle, where we show details in Tab. 1.
Note that both EI and BE focus on fairness across different groups. To comprehensively compare different algorithms, we also evaluate the unfairness induced by each algorithm. Specifically, we consider both fairness notions with and without considering agent behavior, including Equal Improvability (EI), Bounded Effort (BE), Demographic Parity (DP), and Equal Opportunity (EO), which are defined in Tab. 1. We run all experiments on a MacBook Pro with Apple M1 Pro chips, memory of 16GB, and Python 3.9.13. In all experiments of Sec. 6.3, we use the ADAM optimizer. Except ERM, we first perform hyper-parameter selections for all algorithms, e.g., we select $\lambda_{1},\lambda_{2}$ of STWF by grid searching with cross-validation. For STWF, we perform cross-validation on 7 different seeds to select the learning rate from $\{0.001,0.01,0.1\}$ and $\lambda_{1},\lambda_{2}\in\{0,0.5,1,1.5,2\}$. For EI, BE and SAFE, we choose the hyperparameters in the same way as the previous works Guldogan et al. (2022); Rosenfeld et al. (2020); Heidari et al. (2019). Specifically, $\lambda$ (the strength of the regularization) is $0.1$ for all these algorithms.
Table 1: Fairness notions used in our experiments.
Notion Definition
(a) Comparisons under the synthetic dataset.
(b) Comparisons under the German Credit dataset.
(c) Comparisons under the ACSincome-CA dataset.
@@ 6.3 Comparison with Previous Algorithms
$\mathbf{-0.15}\pm 0.00$
6.3.1 Comparison of welfare.

### v2

@@ 3.2 Welfare of different parties
where $P_{XY}$ is the joint distribution of $(X,Y)$. When $u(D(x),y)=\textbf{1}(D(x)=y)$, $\text{DW}(f)$ reduces to the prediction accuracy. In Sec. 5 and 6, we focus on this special case, but the results and analysis can be extended to other utility functions.
@@
$\text{IMP}(f)$ measures the increase in the overall population qualification caused by the agent response. This notion has been theoretically and empirically studied under linear settings (e.g., Kleinberg and Raghavan (2020); Bechavod et al. (2022)).
@@
It illustrates whether there are agents suffering from deteriorating qualifications after they best respond. This notion has been studied in Bechavod et al. (2022) and was interpreted as "do no harm" constraint, i.e., the deployed decisions should avoid adversely impacting agents. Rosenfeld et al. (2020) has shown that agent safety can not be automatically guaranteed even under linear settings, and they proposed an algorithm to ensure safety in general settings. However, they did not consider agent improvement which we believe is an equally important measure of social welfare.
@@
Previous literature rarely discussed the welfare of agents. The closest notion was social burden (Milli et al. 2019) defined as the minimum cost a positively-labeled user must pay to obtain a positive decision, where the authors assume agents can only “game" the decision-maker without improving their true qualifications. Since the agents can always strive to improve in our setting, their efforts are not necessarily a “burden". However, a relatively accurate policy may intentionally underestimate the current qualification to incentivize agent improvement. Thus, we propose agent welfare which measures the extent to which the agent qualification is under-estimated by the decision-maker, i.e.,
@@ 4 Welfare Analysis
This section explores the relationships between the decision-maker, social, and agent welfare introduced in Sec. 3. Among all prior works, (Bechavod et al. 2022) provides a theoretical analysis of social and decision-maker welfare under linear settings. In this section, we first revisit the results in (Bechavod et al. 2022) and show that they may not be applicable to non-linear settings. Then, we examine the relationships between three types of welfare generally.
@@ 4.1 Beyond the linear setting
Consider a hiring scenario where a company needs to assess each applicant’s fitness for a position. Each applicant has a feature $X\in[0,1]$ demonstrating his/her ability (e.g., coding ability for a junior software engineer position). Suppose the labeling function $h(x)=-4x\cdot(x-1)$ is quadratic with maximum taken at $x=0.5$, which implies that an applicant’s fitness will decrease if the ability is too low ("underqualified") or too high ("overqualified")222”Overqualified” means an applicant has skills significantly exceeding the requirement. He/she may fit more senior positions or desire a much higher salary than this position can offer.. Under the quadratic labeling function, if the decision-maker can only select a decision policy $f$ from a linear hypothesis class $\mathcal{F}$ (i.e., violating the requirement that $h\in\mathcal{F}$ in Thm. 4.1), then the optimal policies maximizing decision-maker welfare and agent improvement are completely different, i.e., $\text{DW}(f)$ and $\text{IMP}(f)$ cannot be maximized simultaneously. Moreover, agent safety $\text{SF}(f)$ is also not guaranteed. To illustrate this, consider the example in Fig. 1 with five agents: four green agents with features below $0.5$ and a black one above $0.5$, then the policy that maximizes $\text{IMP}(f)$ can cause the qualification of black agent to deteriorate and lead to negative agent safety $\text{SF}(f)<0$.
@@
Under the same setting of Example 4.2, instead of restricting $f$ to be linear, let the decision-maker deploy $f(x)=h(x)=-4x\cdot(x-1)$. Obviously, $f$ maximizes both the decision-maker welfare and agent welfare. However, it can hurt social welfare. Consider an example shown in Fig. 2 where all agents have the same feature $x=0.4$ and information level $K=1$ (violating the information level requirement of Thm. 4.1). Their cost function is $c(x,x^{\prime})=(x^{\prime}-x)^{2}$. Since $\nabla f(x)=-8x+4=0.8$, agents derive $Q^{f}_{x}(x^{\prime})=f(x)+0.8\cdot(x^{\prime}-x)=0.8x^{\prime}+0.64$. Then agent post-response feature will be $x^{*}=x+\frac{0.8}{2}=0.8$. As a result, the agent qualification becomes $h(x^{*})=h(0.8)=0.64<h(0.4)=0.96$, which is worse than the initial qualification, i.e., agent safety is not optimal by deploying $f=h$.
Figure 2: Violation of agent safety $\text{SF}(f)$ when $f=h$ is quadratic (the blue curve): All agents have the same feature value concentrating on $X=0.4$ (the solid black circle). If applying the ground truth as the decision policy, all agents will respond by moving to $x^{*}=0.8$ (the hollow black circle). Although the policy maximizes $\text{AW}(f)$ and $\text{DW}(f)$, all agents become less qualified after they respond, harming the social welfare. Both $\text{IMP}(f)$ and $\text{SF}(f)$ are smaller than 0.
@@ 4.2 Compatibility analysis of each welfare pair
Decision-maker welfare & agent welfare
@@
Decision-maker welfare & social welfare
Unlike compatibility results in Sec 4.2, the decision-maker welfare and either element of social welfare (i.e., agent improvement $\text{IMP}(f)$ or agent safety $\text{SF}(f)$) in general are not compatible, and their compatibility needs all conditions in Thm. 4.1 to be satisfied.
@@
Agent improvement & agent safety
@@
Suppose agents have information level $K$ and respond to $f$ based on Def. 3.2. Let $^{*}\in\arg\max_{f\in\mathcal{F}}\text{IMP(f)}$(f) be any policy that maximizes the agent improvement, and let $Q^{f^{*}}_{x}$ be the estimation of $f^{*}$ for agents with feature vector $x$. Then any $f^{*}$ is guaranteed to be safe (i.e., $\text{SF}(f^{*})=0$) regardless of the feature distribution $P_{X}$ and the cost function $c(x,x^{\prime})$ if and only if either of the following scenarios holds: (i) Thm. 4.1 holds; (ii) For each agent with feature vector $x$, we have: $\frac{\partial h}{\partial\widetilde{x}_{i}}\cdot\frac{\partial Q^{f^{*}}_{x}}{\partial\widetilde{x}_{i}}\geq 0$ for any $\widetilde{x}=[\widetilde{x}_{1},...,\widetilde{x}_{d}]\in\text{dom}(X)$.
@@
Social welfare & agent welfare
@@
Suppose agents have information level $K$ and respond to $f$ based on Def. 3.2. Denote =$\mathcal{F}^{\prime}=\{f:f\in\mathcal{F}~{}and~{}f(x)\geq h(x),\forall x\in\text{dom}(X)\}$ as the set of policies maximizing the agent welfare. Then for any $f^{*}\in\arg\max_{f\in\mathcal{F}}\text{IMP}(f)$, $f^{*}$ is guaranteed to also maximize $\text{AW}(f)$ regardless of the feature distribution $P_{X}$ and the cost function $c(x,x^{\prime})$ if and only if either of the following scenarios holds: (i) Thm. 4.1 holds; (ii) $\exists f^{\prime}\in\mathcal{F}^{\prime}$ and constant $C\in\operatorname{\mathbb{R}}^{+}$ such that $Q^{f^{\prime}}_{x}=Q^{f^{*}}_{x}+C$ for any $x\in\text{dom}(X)$.
Again, scenario (i) in Thm. 4.7 is a sufficient condition, while scenario (ii) means that the decision-maker can find a classifier that induces perfect agent welfare while simultaneously giving the agents the same information as $f^{*}$. Note that $Q^{f^{\prime}}_{x}=Q^{f^{*}}_{x}+C$ not necessarily implies $f^{\prime}=f^{*}+C$ when $\text{dom}(X)$ is a zero-measure set. For example, when $K=2$ and $\text{dom}(X)=\{0\}$, we can have $f^{*}=\text{e}^{x}-x-1$ and $f^{\prime}=\frac{x^{2}}{2}+0.5$, where $f^{*}$ belongs to the exponential function family but $f^{\prime}$ belongs to the quadratic function family.
To conclude, only decision-maker welfare and agent welfare can be maximized at the same time relatively easily, other pairs of welfare need strong conditions to align under the general settings. Even the two notions in social welfare (i.e., agent improvement and safety) can easily contradict each other, and the decision-maker welfare and social welfare need the most restrictive conditions (which is exactly Thm. 4.1) to be aligned. These results highlight the difficulties of balancing the welfare of different parties in general non-linear settings, which further motivates the "irreducible" optimization framework we will introduce in Sec. 5.
@@ 5 Welfare-Aware Optimization
Algorithm 1 Strategic Welfare-Aware Optimization (STWF)
@@
Output: Classifier $f\in\mathcal{F}$
@@
10:     end for
11:  end for
12:  Return $\theta$
@@
The optimizer of  (1) is Pareto optimal and it can be solved using gradient-based algorithms, as long as the gradient of losses for the model parameter (denoted as $\theta$) exists. Since $f,h$ are continuous, gradients already exist for $\mathbf{\ell_{DW}}$ and $\mathbf{\ell_{AW}}$. For $\mathbf{\ell_{SWF}}$, because it is a function of post-response feature $X^{*}$, we need gradients to exist for $X^{*}$ as well. If we already know how agents respond to the model (i.e., $X^{*}$ as a function of $f$ illustrated in Def. 3.2), then we only need the $K+1^{th}$ gradients to exist for $f(x)$ with respect to $\theta$. When $K=1$, the objective (1) can be easily optimized if $\nabla f(x)$ is continuous in $\theta$, which is a relatively mild requirement and has also been used in previous literature (Rosenfeld et al. 2020). The complete algorithm for StraTegic WelFare-Aware optimization (STWF) is shown in Alg. 1.
@@
6:  end for
7:  Train $\Delta_{\phi}$ with $\mathcal{S}$
8:  Return $\phi$
Alg. 2 requires a population for controlled experiments. As pointed out by Miller, Milli, and Hardt (2020), we can never learn the agent response without the controlled trials. Though the availability of controlled experiments has been assumed in most literature on performative prediction (Perdomo et al. 2020; Izzo, Ying, and Zou 2021), it may be costly to conduct the experiments in practice.
@@ 6 Experiments
$-0.16\pm 0.00$
(a) Welfare while increasing $\lambda_{1}$ and keeping $\lambda_{2}=0$.
(b) Welfare while increasing $\lambda_{2}$ and keeping $\lambda_{1}=0$.
Figure 3: Effects of adjusting $\lambda_{1},\lambda_{2}$ on the welfare of each party: synthetic data (left plot), German Credit data (middle plot), ACSIncome-CA data (right plot). Increasing $\lambda_{1}$ ($\lambda_{2}$) results in improved SWF (AW) and trade-offs exist.
(a) Comparisons under the synthetic dataset.
(b) Comparisons under the German Credit dataset.
(c) Comparisons under the ACSincome-CA dataset.
We conduct comprehensive experiments on one synthetic and two real-world datasets to validate STWF [^3] https://github.com/osu-srml/StrategicWelfare. We first illustrate the trade-offs for each pair of welfare and examine how they can be balanced by adjusting $\lambda_{1},\lambda_{2}$ in Alg. 1. Then, we compare STWF with previous algorithms that only consider the welfare of a subset of parties. Since many of the existing algorithms also consider fairness across agents from different groups, we also evaluate the unfairness induced by our algorithm in addition to decision-maker, agent, and social welfare. We use a three-layer ReLU neural network to learn labeling function $h$ and logistic regression to train the policy $f$, and $D(x)=\mathbf{1}(f(x)\geq 0.5)$. For each parameter setting, we perform 20 independent runs of the experiments with different random seeds and report the average.
6.1 Datasets
  • ACSIncome-CA data (Ding et al. 2021). It contains income-related data for California residents. Let label $Y$ be whether an agent’s annual income is higher than $50$K USD and let the sensitive attribute $Z$ be sex. We randomly select $20000$ samples from the data of year $2018$. We assume all agents have the same effort budget $c(x,x^{\prime})=\|x^{\prime}-x\|^{2}$.
  • German Credit data (Dua and Graff 2017). It has $1000$ samples and the label $Y$ is whether the agent’s credit risk is high or low. We adopt the data preprocessed by (Jiang and Nachum 2020) where age serves as the sensitive attribute to split agents into two groups. We assume the agents are able to improve the following features: existing account status, credit history, credit amount, saving account, present employment, installment rate, guarantors, residence. We assume all agents have the same effort budget $c(x,x^{\prime})=\|x^{\prime}-x\|^{2}$.
6.2 Balance the Welfare with Different $\lambda_{1},\lambda_{2}$
We first validate STWF and show that adjusting the hyper-parameters $\lambda_{1},\lambda_{2}$ effectively balances the welfare of different parties. The results in Fig. 3 are as expected, where social welfare $\text{SWF}(f)$ increases in $\lambda_{1}$ and agent welfare $\text{AW}(f)$ increases in $\lambda_{2}$. This shows the effectiveness of using our algorithm to adjust the allocation of welfare under different situations. As we analyzed in Sec. 4, under the general non-linear setting, each pair of welfare possibly conflicts. Fig. 3 validates this by demonstrating the trade-offs between different welfare pairs on two datasets, while it also reveals that the conflicts are not always serious between each pair. For the synthetic data, increasing social welfare slightly conflicts with decision-maker welfare, while increasing agent welfare drastically conflicts with the decision-maker welfare. But the trade-off between social welfare and agent welfare is not obvious. For the German Credit dataset, increasing social welfare seems to slightly conflict with both agent welfare and decision-maker welfare, while increasing agent welfare has an obvious trade-off with decision-maker welfare. Conversely, for the ACSIncome-CA dataset, increasing social welfare only sacrifices agent welfare, but increasing agent welfare incurs conflicts with both other welfare.
6.3 Comparison with Previous Algorithms
Specifically, we compare STWF with four algorithms:
  • Safety (SAFE) (Rosenfeld et al. 2020): It considers both decision-maker welfare and agent safety. The goal is to train an accurate model without hurting agent qualification.
  • Equal Improvement (EI) (Guldogan et al. 2022): It considers agent improvement by equalizing the probability of unqualified agents becoming qualified from different groups, as opposed to maximizing agent improvement considered in our work.
  • Bounded Effort (BE) (Heidari, Nanda, and Gummadi 2019): It considers agent improvement by equalizing the proportion of new agents becoming qualified after the best response. The difference between BE and EI is a bit subtle, where we show details in Tab. 2.
Note that both EI and BE focus on fairness across different groups. To comprehensively compare different algorithms, we also evaluate the unfairness induced by each algorithm. Specifically, we consider both fairness notions with and without considering agent behavior, including Equal Improvability (EI), Bounded Effort (BE), Demographic Parity (DP), and Equal Opportunity (EO), which are defined in Tab. 2. We run all experiments on a MacBook Pro with Apple M1 Pro chips, memory of 16GB, and Python 3.9.13. In all experiments of Sec. 6.3, we use the ADAM optimizer. Except ERM, we first perform hyper-parameter selections for all algorithms, e.g., we select $\lambda_{1},\lambda_{2}$ of STWF by grid searching with cross-validation. For STWF, we perform cross-validation on 7 different seeds to select the learning rate from $\{0.001,0.01,0.1\}$ and $\lambda_{1},\lambda_{2}\in\{0,0.5,1,1.5,2\}$. For EI, BE and SAFE, we choose the hyperparameters in the same way as the previous works (Guldogan et al. 2022; Rosenfeld et al. 2020; Heidari, Nanda, and Gummadi 2019). Specifically, $\lambda$ (the strength of the regularization) is $0.1$ for all these algorithms.
Table 2: Fairness notions used in our experiments.
Notion Definition
Comparison of welfare.